A triangle has one angle that measures 5 degrees more than twice the smallest angle, and the largest angle measures 11 degrees less than 3 times the measure of the smallest angle. Find the measure of all three angles.
Answer by rfer(16322)

x+2x+5+3x-11=180
6x=186
x=31
2x+5=67
3x-11=82